# The Art of Storytelling: Multi-Agent Generative AI for Dynamic Multimodal Narratives

### v4

@@ Abstract
This paper introduces the concept of an education tool that utilizes Generative Artificial Intelligence (GenAI) to enhance storytelling for children. The system combines GenAI-driven narrative co-creation, text-to-speech conversion, text-to-music and text-to-video generation to produce an engaging experience for learners. We describe the co-creation process, the adaptation of narratives into spoken words using text-to-speech models, and the transformation of these narratives into contextually relevant visuals through text-to-video technology. Our evaluation covers the linguistics of the generated stories, the text-to-speech conversion quality, and the accuracy of the generated visuals.
The Art of Storytelling: Multi-Agent Generative AI for
Dynamic Multimodal Narratives
@@ 1 Introduction
In this paper, we introduce a multi-agent system designed to co-create multimodal stories for children. The system is structured around two well-established storytelling frameworks: Freytag’s Pyramid from his book The technique of Drama [^1] https://www.gutenberg.org/files/50616/50616-h/50616-h.htm and Propp’s 31 narrative functions from his book Morphology of the Folktale [^2] https://www.jstor.org/stable/10.7560/783911. Freytag’s Pyramid provides a foundational structure, dividing stories into five phases—exposition, rising action, climax, falling action, and resolution—ensuring that the generated narratives follow a coherent and engaging progression. Propp’s functions serve as a guide for generating the essential elements found in folktales, offering a consistent framework for creating dynamic and imaginative stories that resonate with young audiences. These frameworks can also effectively enhance storytelling skills Ciğerci and Yıldırım (2023). From a pedagogical perspective, these frameworks allow children to actively participate in the storytelling process rather than passively consuming content.
@@
    System Development: We present a novel multi-agent system that integrates various generative AI models, including TTS models to convert generated stories into emotionally expressive narration, TTV models to generate animated visuals that align with the story structure, and a TTM model to create background music that enhances the emotional tone of the narrative.
@@ 4.1.1 LLMs for Story Generation
In order to understand how they evaluate stories submitted by their students, we interviewed teachers who teach English as a subject to students of age group 6 to 12. The teachers emphasized assessing the content of the story for its quality. This includes evaluating the language used, checking the logical sequence of the story, usage of “key terms” to “cue the reader to the direction of the discourse”. The teachers noted:
“…..we try not to use any harmful/abusive language”
“We check if there is a clear beginning and end”
“Does the writer stick to the topic of the story?”
Other than the content, the structure of the story is also scrutinized. One of the teachers said:
“A group of words does not make a complete sentence. Sentences that are made up of a group of words without a direction need to be penalized”

### v5

# The Art of Storytelling: Evaluating AI Models for Storytelling

@@ Abstract
This paper introduces the concept of an education tool that utilizes Generative Artificial Intelligence (GenAI) to enhance storytelling. We evaluate GenAI-driven narrative co-creation, text-to-speech conversion, text-to-music and text-to-video generation to produce an engaging experience for learners. We describe the co-creation process, the adaptation of narratives into spoken words using text-to-speech models, and the transformation of these narratives into contextually relevant visuals through text-to-video technology. Our evaluation covers the linguistics of the generated stories, the text-to-speech conversion quality, and the accuracy of the generated visuals.
The Art of Storytelling: Evaluating AI Models for Storytelling
@@ 1 Introduction
In this paper, we evaluate a the feasibility of GenAI to co-create multimodal stories for children. The system is structured around two well-established storytelling frameworks: Freytag’s Pyramid from his book The technique of Drama [^1] https://www.gutenberg.org/files/50616/50616-h/50616-h.htm and Propp’s 31 narrative functions from his book Morphology of the Folktale [^2] https://www.jstor.org/stable/10.7560/783911. Freytag’s Pyramid provides a foundational structure, dividing stories into five phases—exposition, rising action, climax, falling action, and resolution—ensuring that the generated narratives follow a coherent and engaging progression. Propp’s functions serve as a guide for generating the essential elements found in folktales, offering a consistent framework for creating dynamic and imaginative stories that resonate with young audiences. These frameworks can also effectively enhance storytelling skills Ciğerci and Yıldırım (2023). From a pedagogical perspective, these frameworks allow children to actively participate in the storytelling process rather than passively consuming content.